# Discrete State Diffusion Models: A Sample Complexity Perspective

### v1

@@ 7 Conclusion
We investigate the sample complexity of training discrete-state diffusion models using sufficiently expressive neural networks. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-2})$, which, to our knowledge, is the first such result under this setting. Notably, our analysis avoids exponential dependence on the data dimension and network parameters. For comparison, existing works assume a bound over the efficiency of the trained network, while we have decomposed this error, using SGD for training through a dataset of finite size.
@@ C.2 Bounding the Statistical Error
$\mathcal{E}^{\mathrm{stat}}_{k}\leq\mathcal{O}\left((S-1)d\cdot\sqrt{\frac{{{\log}\left(\frac{2}{\gamma}\right)}}{{n_{k}}}}\right)$ (64)
@@
$\mathbb{E}_{\sigma}\max_{\theta\in\Theta^{{}^{\prime\prime}}}\sum_{i=1}^{n_{k}}f(\theta){\sigma_{i}}\leq||f(\theta)||_{2}\leq(BW)^{L}\!\left((S-1)d.+\frac{L}{W}\right)$ (71)
@@
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{k}(\theta)\right|\right]\leq\mathcal{O}\left(\frac{(W)^{L}\left((S-1)d.+\frac{L}{W}\right)}{n_{k}}\right)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\gamma}}{n_{k}}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}.$ (72)
@@
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{k}(\theta)\right|\right]\leq\mathcal{O}\left((W)^{L}\left((S-1)d.+\frac{L}{W}\right)\cdot\sqrt{\frac{\log\frac{1}{\gamma}}{n_{k}}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}$ (73)
@@
$\|s^{a}_{\theta,k}(x_{k})-s^{b}_{\theta,k}(x_{k})\|^{2}$ $\leq L^{\prime}.\|\theta^{a}_{k}-\theta^{b}_{k}\|^{2}$ (75)
$\leq L^{\prime}.\mu\left|\mathcal{L}_{k}(\theta^{a}_{k})-\mathcal{L}_{k}(\theta^{b}_{k})\right|$ (76)
$\leq\mathcal{O}\left((W)^{L}\left((S-1)d.+\frac{L}{W}\right)\sqrt{\frac{\log\frac{1}{\gamma}}{n_{k}}}\right).$ (77)
@@ C.3 Bounding the Optimization Error
$\mathcal{E}^{\mathrm{opt}}_{k}\leq\mathcal{O}\left((S-1)d.\sqrt{\frac{\log\left(\frac{2}{\gamma}\right)}{n_{k}}}\right).$ (78)
@@
$\left\|\overline{s}_{k}(\mathbf{x}_{k})-s_{k}^{a}(\mathbf{x}_{k})\right\|^{2}$ $\leq L\cdot\left\|\overline{\theta}_{k}-\theta_{k}^{a}\right\|^{2}\leq\left|\mathcal{L}_{k}(\theta_{k})-\mathcal{L}_{k}^{*}\right|$ (93)
@@
$\left\|s_{k}^{a}(\mathbf{x}_{k})-s_{k}^{b}(\mathbf{x}_{k})\right\|^{2}$ $\leq L\cdot\left\|\theta^{a}_{k}-\theta^{b}_{k}\right\|^{2}$ (95)
$\leq L\cdot\mu\left|\mathcal{L}_{k}(\theta^{a}_{k})-\mathcal{L}_{k}(\theta^{b}_{k})\right|$ (96)
$\leq\mathcal{O}\left((W)^{L}\left((S-1)d.+\frac{L}{W}\right)\cdot\sqrt{\frac{\log\frac{2}{\gamma}}{n_{k}}}\right)$ (97)
@@
$\leq\mathcal{O}\left(\log\left(\frac{1}{n}\right)\right)+\mathcal{O}\left((W)^{L}\left((S-1)d.+\frac{L}{W}\right)\cdot\sqrt{\frac{\log\frac{2}{\gamma}}{n_{k}}}\right)$
$\leq\mathcal{O}\left((W)^{L}\left((S-1)d.+\frac{L}{W}\right)\cdot\sqrt{\frac{\log\frac{2}{\gamma}}{n_{k}}}\right)$ (98)
@@ D.2 KL divergence bound
We shall now focus on bounding the third term in (107) which expresses the score approximation error by writing $A_{k}=\mathbb{E}_{\mathbf{x}_{(k+1)h}\sim q_{(k+1)h}}\left\|s_{(k+1)h}(\mathbf{x}_{(k+1)h})-\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{(k+1)h})\right\|_{2}^{2}$. For notational simplicity we re-write $A_{k}=\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\left\|s_{k}(\mathbf{x}_{k})\right.-\left.\hat{s}_{\theta,k}(\mathbf{x}_{k})\right\|_{2}^{2}$ as mentioned in (27). For $D_{\mathrm{KL}}(q_{0}\|p_{T})\leq\tilde{O}(\epsilon)$ to hold, we have to ensure that each of the terms from (107) scales as $\tilde{O}(\epsilon)$. Therefore, we bound the score-approximation error term $\frac{C^{3}}{S}\sum_{k=0}^{K-1}h\,\mathbb{E}_{\mathbf{x}_{(k+1)h}\sim q_{(k+1)h}}\bigl\|s_{(k+1)h}(\mathbf{x}_{(k+1)h})-\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{(k+1)h})\bigr\|_{2}^{2}$ based on the number of samples $n_{k}$ required to achieve an $\epsilon_{score}$-accurate score estimator.
@@ Appendix E Hardness of Learning in Discrete-State Diffusion Models
Although $P$ and $Q$ are separated by $D_{\mathrm{KL}}(P\|Q)>20\varepsilon$, their means differ only by $\Theta(\varepsilon)$, i.e., their separation is linear in $\varepsilon$. If a simulator $R$ achieves $D_{\mathrm{KL}}(U\|R)\leq\varepsilon$ for some $U\in\{P,Q\}$, it cannot simultaneously satisfy this bound for both distributions. Therefore, distinguishing whether the samples originated from $P$ or $Q$ means that we can distinguish two distributions whose means differ by $\Theta(\varepsilon)$. By standard mean-testing lower bounds (Wainwright [2019]), distinguishing two distributions whose means differ by $\Theta(\varepsilon)$ requires at least $\Omega(1/\varepsilon^{2})$ samples. Hence, any learner that produces such an $R$ with $D_{\mathrm{KL}}(U\|R)\leq\varepsilon$ must necessarily use $\Omega(1/\varepsilon^{2})$ samples.

### v2

@@ 7 Conclusion
We investigate the sample complexity of training discrete-state diffusion models using sufficiently expressive neural networks. We derive a sample complexity bound of $\widetilde{\mathcal{O}}(\epsilon^{-2})$, which, to our knowledge, is the first such result under this setting. Notably, our analysis avoids exponential dependence on the data dimension. For comparison, existing works assume a bound over the efficiency of the trained network, while we have decomposed this error, using SGD for training through a dataset of finite size.
Recent work by Sahoo et al. (2026) shows that rectified flow matching (RFM) also achieves an $\widetilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity, improving upon the $\widetilde{\mathcal{O}}(\epsilon^{-4})$ rate known for continuous flow matching (Gaur et al., 2025a). Developing guarantees for discrete flow matching counterparts remains an important direction for future work, with the potential to improve practical performance. Additionally, a more refined treatment of approximation error arising from finite-capacity neural networks remains an open and promising avenue for further study.
@@ C.2 Bounding the Statistical Error
$\mathcal{E}^{\mathrm{stat}}_{k}\leq\mathcal{O}\left((S-1)d\sqrt{\frac{{{\log}\left(\frac{2}{\gamma}\right)}}{{n_{k}}}}\right)$ (62)
@@
[…]
@@
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{k}(\theta)\right|\right]\leq\mathcal{O}\left(\frac{(W)^{L}\left((S-1)d+\frac{L}{W}\right)}{n_{k}}\right)+\mathcal{O}\left(\sqrt{\frac{\log\frac{1}{\gamma}}{n_{k}}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}.$ (70)
@@
$\mathbb{E}\left[\left|\mathcal{L}_{k}(\theta)-\widehat{\mathcal{L}}_{k}(\theta)\right|\right]\leq\mathcal{O}\left((W)^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\frac{\log\frac{1}{\gamma}}{n_{k}}}\right),~~\quad\forall~\theta\in\Theta^{\prime\prime}$ (71)
@@
$\|s^{a}_{\theta,k}(x_{k})-s^{b}_{\theta,k}(x_{k})\|^{2}$ $\leq L^{\prime}\|\theta^{a}_{k}-\theta^{b}_{k}\|^{2}$ (73)
$\leq L^{\prime}\mu\left|\mathcal{L}_{k}(\theta^{a}_{k})-\mathcal{L}_{k}(\theta^{b}_{k})\right|$ (74)
$\leq\mathcal{O}\left((W)^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\frac{\log\frac{1}{\gamma}}{n_{k}}}\right)$ (75)
@@ C.3 Bounding the Optimization Error
$\mathcal{E}^{\mathrm{opt}}_{k}\leq\mathcal{O}\left((S-1)d\sqrt{\frac{\log\left(\frac{2}{\gamma}\right)}{n_{k}}}\right).$ (76)
@@
$\left\|\overline{s}_{k}(\mathbf{x}_{k})-s_{k}^{a}(\mathbf{x}_{k})\right\|^{2}$ $\leq L\left\|\overline{\theta}_{k}-\theta_{k}^{a}\right\|^{2}\leq\left|\mathcal{L}_{k}(\theta_{k})-\mathcal{L}_{k}^{*}\right|$ (91)
@@
$\left\|s_{k}^{a}(\mathbf{x}_{k})-s_{k}^{b}(\mathbf{x}_{k})\right\|^{2}$ $\leq L\left\|\theta^{a}_{k}-\theta^{b}_{k}\right\|^{2}$ (93)
$\leq L\mu\left|\mathcal{L}_{k}(\theta^{a}_{k})-\mathcal{L}_{k}(\theta^{b}_{k})\right|$ (94)
$\leq\mathcal{O}\left((W)^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\frac{\log\frac{2}{\gamma}}{n_{k}}}\right)$ (95)
@@
$\leq\mathcal{O}\left(\log\left(\frac{1}{n}\right)\right)+\mathcal{O}\left((W)^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\frac{\log\frac{2}{\gamma}}{n_{k}}}\right)$
$\leq\mathcal{O}\left((W)^{L}\left((S-1)d+\frac{L}{W}\right)\sqrt{\frac{\log\frac{2}{\gamma}}{n_{k}}}\right)$ (96)
@@ D.2 KL divergence bound
We shall now focus on bounding the third term in (105) which expresses the score approximation error by writing $A_{k}=\mathbb{E}_{\mathbf{x}_{(k+1)h}\sim q_{(k+1)h}}\left\|s_{(k+1)h}(\mathbf{x}_{(k+1)h})-\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{(k+1)h})\right\|_{2}^{2}$. For notational simplicity we re-write $A_{k}=\mathbb{E}_{\mathbf{x}_{k}\sim q_{k}}\left\|s_{k}(\mathbf{x}_{k})\right.-\left.\hat{s}_{\theta,k}(\mathbf{x}_{k})\right\|_{2}^{2}$ as mentioned above (28). For $D_{\mathrm{KL}}(q_{0}\|p_{T})\leq\tilde{O}(\epsilon)$ to hold, we have to ensure that each of the terms from (105) scales as $\tilde{O}(\epsilon)$. Therefore, we bound the score-approximation error term $\frac{C^{3}}{S}\sum_{k=0}^{K-1}h\,\mathbb{E}_{\mathbf{x}_{(k+1)h}\sim q_{(k+1)h}}\bigl\|s_{(k+1)h}(\mathbf{x}_{(k+1)h})-\hat{s}_{\theta,(k+1)h}(\mathbf{x}_{(k+1)h})\bigr\|_{2}^{2}$ based on the number of samples $n_{k}$ required to achieve an $\epsilon_{score}$-accurate score estimator.
@@ Appendix E Hardness of Learning in Discrete-State Diffusion Models
Although $P$ and $Q$ are separated by $D_{\mathrm{KL}}(P\|Q)>20\varepsilon$, their means differ only by $\Theta(\varepsilon)$, i.e., their separation is linear in $\varepsilon$. If a simulator $R$ achieves $D_{\mathrm{KL}}(U\|R)\leq\varepsilon$ for some $U\in\{P,Q\}$, it cannot simultaneously satisfy this bound for both distributions. Therefore, distinguishing whether the samples originated from $P$ or $Q$ means that we can distinguish two distributions whose means differ by $\Theta(\varepsilon)$. By standard mean-testing lower bounds [Wainwright, 2019], distinguishing two distributions whose means differ by $\Theta(\varepsilon)$ requires at least $\Omega(1/\varepsilon^{2})$ samples. Hence, any learner that produces such an $R$ with $D_{\mathrm{KL}}(U\|R)\leq\varepsilon$ must necessarily use $\Omega(1/\varepsilon^{2})$ samples.
Appendix F Dimension of the discrete-state score
In principle, if we view the discrete-state diffusion process as a continuous-time Markov chain (CTMC) on the full state space $\mathcal{X}=[S]^{d}$ with $N=S^{d}$ states, then the forward rate matrix $Q_{t}$ would be of size $N\times N$, and the corresponding discrete-state score $s_{t}(\mathbf{x})=\bigl(q_{t}(\mathbf{y})/q_{t}(\mathbf{x})\bigr)_{\mathbf{y}\neq\mathbf{x}}$ would indeed live in $\mathbb{R}^{N-1}$. However, such a formulation is computationally intractable for high-dimensional data.
Following prior works [Chen and Ying, 2024] and [Zhang et al., 2025], we impose a coordinate-wise (token-wise) factorization of the forward process. Specifically, each of the $d$ coordinates evolves independently according to a shared $S\times S$ rate matrix. Under this assumption, every jump of the forward CTMC modifies exactly one coordinate at a time, and transitions are only allowed between states whose Hamming distance is one. As a result, the full rate matrix $Q_{t}$ is extremely sparse: from any state $\mathbf{x}\in[S]^{d}$, there are only d(S1)d(S-1) nonzero off-diagonal transitions.
The reverse-time CTMC inherits exactly the same sparsity structure. Its rate matrix $Q_{t}^{\leftarrow}$ involves only transitions of the form
$\mathbf{x}\;\longrightarrow\;\mathbf{x}_{\setminus i}\odot\hat{x}_{i},\qquad i\in[d],\ \hat{x}_{i}\neq x_{i},$
and for each such transition the reverse rate is given by
$Q_{t}^{\leftarrow}\!\left(\mathbf{x},\mathbf{x}_{\setminus i}\odot\hat{x}_{i}\right)\;=\;Q^{\mathrm{tok}}_{T-t}(\hat{x}_{i},x_{i})\,\frac{q_{t}(\mathbf{x}_{\setminus i}\odot\hat{x}_{i})}{q_{t}(\mathbf{x})}.$
Thus, the only probability ratios that appear in the reverse dynamics are those corresponding to single-coordinate replacements.
This observation leads us to express the discrete-state score in the form required by the reverse-time CTMC. We therefore define
$s_{t}(\mathbf{x})_{i,\hat{x}_{i}}\;:=\;\frac{q_{t}(\mathbf{x}_{\setminus i}\odot\hat{x}_{i})}{q_{t}(\mathbf{x})},\qquad i\in[d],\ \hat{x}_{i}\neq x_{i},$
which parametrizes all nonzero transition rates of $Q_{t}^{\leftarrow}$. For each coordinate $i$, the collection $\{s_{t}(\mathbf{x})_{i,\hat{x}_{i}}:\hat{x}_{i}\neq x_{i}\}$ has S1S-1 degrees of freedom, since the probabilities over the $S$ possible symbols at that coordinate must sum to one after normalization. Aggregating across all $d$ coordinates yields a score representation in $\mathbb{R}^{d(S-1)}$.
Importantly, this reduction from $\mathbb{R}^{S^{d}-1}$ to $\mathbb{R}^{d(S-1)}$ is not an approximation or modeling choice. It is a consequence of the sparsity of the forward and reverse generators under the factorized CTMC assumption, and it provides a complete and tractable characterization of the reverse diffusion dynamics.
Appendix G PL Condition and Sample-complexity Scaling
This section clarifies the role of the Polyak–Łojasiewicz (PL) condition in our analysis and explains how the stated sample-complexity bounds should be interpreted.
G.1 Role and Interpretation of the PL Condition
This section clarifies the role of the Polyak–Łojasiewicz (PL) condition in the analysis and explains how it should be interpreted within the theoretical framework of our work.
The PL condition is imposed as a structural regularity assumption on the population loss landscape $\mathcal{L}_{k}(\theta)$ at each discrete time index $k$. Its purpose is to enable control of the optimization error by relating the gradient norm to function-value suboptimality. Importantly, this assumption is not tied to strong convexity, nor is it intended to characterize the degree of overparameterization of the neural network used to parameterize the score.
While overparameterized neural networks form one prominent class of models where PL-type or gradient-dominance conditions are known to hold, such conditions have been established in substantially broader settings. In particular, PL or closely related gradient-dominance properties are known to hold even in regimes that are not heavily overparameterized, i.e., when the number of parameters doesn’t exceed the number of effective constraints.
Prior work by Zhou and Liang [2017] rigorously characterizes gradient-dominance and regularity conditions for several nonconvex architectures, including deep linear networks, deep linear residual networks, and one-hidden-layer nonlinear networks, in neighborhoods of global minimizers, without requiring a heavily overparameterized regime. These results show that even relatively modestly sized networks can exhibit a PL-type landscape, at least locally around solutions. This type of geometric behavior is precisely what is required in our analysis: a condition ensuring that small gradient norms imply small suboptimality, which in turn yields linear-type convergence rates for stochastic gradient methods.
In our framework, the PL condition is imposed on each time-indexed population loss $\mathcal{L}_{k}(\theta)$, rather than on the empirical objective. It is used solely as an analytical device to convert bounds on gradient norms into bounds on function-value suboptimality, thereby enabling control of both statistical and optimization errors. The sample-complexity requirement appearing in the main results should therefore be interpreted as a sufficient condition ensuring that these errors are controlled at the desired level. This requirement does not preclude the population loss from satisfying a PL condition in regimes where the number of samples is comparable to, or larger than, the number of trainable parameters.
G.2 Sample-Complexity Scaling and Width Assumptions
This section clarifies the interpretation of the sample-complexity bounds derived in the paper and explains the role of the width condition appearing in Lemma 1.
The main analysis yields a sample-complexity bound of the form
$n\gtrsim\Omega\!\left(W^{2L}\right),$ (133)
where $W$ denotes the width of the neural network parameterizing the score function and $L$ denotes its depth. This dependence arises from standard Rademacher complexity arguments for deep neural networks, which are known to be conservative, particularly with respect to depth. As a result, the bound should be interpreted as a worst-case sufficient condition ensuring convergence, rather than as a sharp or practically tuned estimate of the number of training samples required. Exponential or highly pessimistic depth dependence is common in current deep-learning theory such as in Benton et al. [2024b] and often substantially overestimates actual data requirements.
When this bound is combined with the sufficient-width condition $W\geq(S-1)d$ used in Lemma 1, the resulting scaling becomes $\Omega((Sd)^{2L})$, which can be numerically large for large vocabularies $S$ or deep networks $L$. This behavior reflects the proof technique and worst-case generalization guarantees rather than an inherent limitation of discrete-state diffusion models. Importantly, the substitution W=(S1)dW=(S-1)d is not required by the main generalization theorem itself; it is introduced solely as a convenient sufficient condition in Lemma 1 to ensure zero approximation error via an explicit construction.
More generally, when W<(S1)dW<(S-1)d such as with parameter sharing or low-rank structure, the analysis extends naturally by introducing an explicit approximation error term. Let
$\varepsilon_{\mathrm{approx}}(W,S,d)$ (134)
denote the infimum over all width-$W$ networks in the considered class of the population approximation error to the target score function. Under this formulation, Lemma 1 and the main excess-risk bound can be restated with an additional additive term $\varepsilon_{\mathrm{approx}}(W,S,d)$. In the special case $W\geq(S-1)d$, the explicit construction implies $\varepsilon_{\mathrm{approx}}(W,S,d)=0$, recovering the original statement of Lemma 1.
This perspective is standard in learning theory. Approximation error induced by restricting to a given function class is often treated as a fixed constant once the class is chosen, allowing the analysis to focus on estimation and optimization errors that depend on the sample size. For example, in PAC-Bayesian analyses, approximation or misspecification errors are absorbed into a constant term once the hypothesis class is fixed [Mai, 2025]. Similarly, in NTK or RKHS based analyses of neural networks, it is typically assumed that the target function lies in, or is well approximated by, the chosen function class, with the resulting misspecification error represented as a constant additive term [Bing et al., 2025]. Existing theoretical analyses of diffusion models adopt analogous assumptions, treating approximation error as fixed [Gupta et al., 2024, Guan et al., 2025, Gaur et al., 2025a].
Appendix H Relationship to Prior Work
This work is inspired by existing theoretical analyses of discrete-state diffusion models [Zhang et al., 2025, Chen and Ying, 2024], particularly those that decompose the overall error into interpretable components and study convergence under idealized assumptions. Similar decomposition is used in works in reinforcement learning such as Gaur et al. [2023a, b, 2024]. We adopt a similar high-level decomposition as a conceptual starting point, but develop technical tools and results tailored to discrete-state diffusion models trained with finite-width neural networks under nonconvex objectives.
At a high level, the novelty of this work lies in making the abstract components of prior discrete-state diffusion analyses explicit, data-dependent, and architecture-aware in the discrete-state setting.
Approximation error (Lemma 1).
Prior analyses typically assume realizability of the score function or work in a fixed functional class, thereby avoiding explicit treatment of approximation error. In contrast, Lemma 1 provides a discrete-state-specific approximation result that characterizes when a finite-width neural network can approximate the discrete-state score function and how this depends on the vocabulary size $S$ and dimension $d$. This lemma explicitly connects neural network expressivity to the statistical behavior of discrete-state diffusion models, a connection that does not appear in existing analyses.
Statistical error (Lemma 2).
While previous works derive convergence bounds at an abstract functional level, our analysis yields an explicit, training data dependent sample-complexity bound under a nonconvex PL-type landscape. This result makes the residual error term in prior analyses concrete and interpretable in terms of network width, depth, vocabulary size, and diffusion time. As a result, the statistical component of the error decomposition becomes an explicit quantity rather than an implicit asymptotic remainder.
Optimization error (Lemma 3).
We further analyze how optimization error propagates through the diffusion time horizon. This discrete-time stability analysis quantifies how local errors accumulate across steps and affect the final generated distribution. Such propagation effects are absent from prior analyses as they assume access to an empirical risk minimizer (ERM) and are essential for understanding convergence in the discrete-state setting with finite optimization steps.
Together, these results extend existing theory to a substantially more realistic setting, yielding guarantees that are not accessible through existing analyses.